OpenStudy (vortish):

Use the multiplication property of inequality to graph the solution set 1/9t is greater or = 6

OpenStudy (anonymous):

\[\frac{ 1 }{ 9t }\ge6 \]

OpenStudy (vortish):

yes that is correct

OpenStudy (anonymous):

cross multiply

OpenStudy (vortish):

6/54t ?

OpenStudy (anonymous):

1/9t is greater or = to 6/1 and cross multiply

OpenStudy (vortish):

\[\frac{ 1 }{ 9 } t \ge 6\]

OpenStudy (vortish):

54/1 t

OpenStudy (anonymous):

multiply 9t by 6

OpenStudy (vortish):

then you get 54t

OpenStudy (anonymous):

yep since u are cross multiplying \[1\ge54t\]

OpenStudy (anonymous):

u then simplify that

OpenStudy (vortish):

ok moyo so that is the solution set for this problem how do you simplfie that

OpenStudy (anonymous):

ivide through by 54 to get \[\frac{ 1 }{ 54}\ge t\]d

OpenStudy (anonymous):

Thanks Vortish